answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How do you get if you roll a die 100 times how many times would you expect to get 1?

If you roll a die 100 times, you would expect to get a 1 about 17 times, because the probability of getting a 1 is 1 in 6, or 0.1667. However, that is theoretical probability; experimental probability - the actual results of doing this 100 times - might not be 17, but if you did this a large number of times, the experimental results would indeed begin to approach the theoretical results.

What is the probability if rolling a sum of 7 or 8 given the second die is a 1 or a 2?

1/4 = 0.25

If the second die is 1 or 2, then there are 12 possible outcomes.

To score 7: 5+2 or 6+1 ⇒ 2 rolls

To score 8: 6+2 ⇒ 1 roll

⇒ to score 7 or 8: 2 + 1 = 3 rolls

So there are 3 rolls to score the required sums out of 12 possible rolls, giving a probability of:

3/12 = 1/4 = 0.25

What is the difference of a normal distribution and a stardard normal distribution?

The standard normal distribution is a special case of the normal distribution. The standard normal has mean 0 and variance 1.

What is the probability of rolling a 7 on a regular six-sided die?

The probability of rolling a 7 on a standard die is 0. It will not ever happen.

How many permutations of the letters in the word Louisiana are there?

The number of permutations of the letters in the word LOUISIANA is 9 factorial or 362,880. However, since the letters I and A are each repeated once, you need to divide that by 4 to determine the number of distinct permutations, giving you 90,720.

What is the z score of x equals 108?

You need the mean and standard deviation in order to calculate the z-score. Neither are given.

What is the probability of a woman having four boys?

-- Zero if she has 0, 1, 2, or 3 children.

-- Roughly 1/16 = 6.25 percent if she has 4 children.

-- Roughly 5/32 = 15.6 percent if she has 5 children.

-- Roughly 15/64 = 23.4 percent if she has 6 children.

What is the probability of rolling a sum less than 6 with two dice?

There are 36 permutations of two dice. Of these, 10 sum to less than six; 1+1, 1+2, 1+3, 1+4, 2+1, 2+2, 2+3, 3+1, 3+2, and 4+1. The probability, then, of rolling a sum less than six is 10 in 36, or 5 in 18, or about 0.2778.

What are the probability of having a male baby in the womb?

Approx 1/2, except that while it is in the womb, it is not a baby!

What is the probability of guess a two digit number?

On a number line, the digits from 10 to 99 are all two digit numbers. How many numeber are that? There is a short cut. You can subtract 10 from 99 and that gives you the number of numbers between 10 and 99 including 10, but you need to add one more to include 99 as well. So, 10 from 99 is 89 adding 1 is 90 ,which is the amount of numbers total that are two digit. The probability of guessing a two digit number is 1/90 =D

How many different ways are there to to scramble to letters of the word mass if s needs to be the first letter?

The number of permutations of the letters MASS where S needs to be the first letter is the same as the number of permutations of the letters MAS, which is 3 factorial, or 6. SMAS SMSA SAMS SASM SSMA SSAM

How do you know if a z score is positive or negative?

A negative Z-Score corresponds to a negative standard deviation, i.e. an observation that is less than the mean, when the standard deviation is normalized so that the standard deviation is zero when the mean is zero.

What is the probability that one randomly-selected person in America will know another randomly-selected person in America?

There is not enough information about the the distribution of the number of people known by each individual - nor the averages. It is therefore no possible to give an answer any more precise than "the probability will be infinitesimally small".

How do you reduce the terms if the probability of drawing a face card is 12 in 52?

You reduce the terms the same way you reduce an ordinary fraction, by finding the greatest common divisor and dividing both numbers by that divisor. In this case, the greatest common divisor is 4, so you divide by 12 and 52 by 4, giving you 3 and 13, so if the probability of drawing a face card is 12 in 52, it is also 3 in 13.

What is the probability that two cards in a 52 card deck are the same color?

In a standard deck the probability is 0. There are not two, but 26 cards of the same colour.

What is the probability of picking 2 hearts at random from a deck of playing cards?

Chance of picking first heart card is 13/52, picking second right after is 12/51, total chance is 13/52 multiplied by 12/51. This works if we pick one card, don't replace it and then pick another. Here is the probability of 2 Hearts if we pick two cards at the same time. 13 Hearts choose 2/ 52 cards choose 2= 13C2/52C2=78/1326=1/17